A nuclear particle was observed to break into two fragments which moved in opposite direction the velocity of each fragment was found to be 2.5×10^8 m/s. relative to the laboratory the velocity of the one fragment relative to other is
1
Expert's answer
2019-05-03T11:31:15-0400

Relative to one of the fragments, the laboratory moves with velocity "v = 2.5 \times 10^8\, \text{m/s}", and, relative to the laboratory, another fragment moves with the same velocity in the same direction. Then the velocity "u" of the second fragment relative to the first one is found by the formula of relativistic addition of velocities:

"u = \frac{v + v}{1 + v \cdot v / c^2} = \frac{2 v}{1 + v^2/c^2} \, ,"


where "c = 3 \times 10^8\, \text{m/s}" is the speed of light. Substituting the numbers, we obtain "u = 2.95 \times 10^8\, \text{m/s}".


Answer: "2.95 \times 10^8\, \text{m/s}".